OpenStudy (anonymous):

Can you check my answers. Please and thank you! Simplify the following expression. 8x - 8y + 5z - 3x - 14y + 3z 11x - 22y + 8z 5x - 6y + 8z 5x - 22y + 8z 11x - 6y + 8z I chose C. Evaluate: -8 - (-3) -11 -5 5 24 I chose C. What is the slope and y-intercept of the graph of 2x + 3y = 18? slope = -2; y-intercept = (0, 18) slope = -2/3; y-intercept = (0, 6) slope = -3/2; y-intercept = (0, 2) slope = 6; y-intercept = (0, ) I chose B.

OpenStudy (anonymous):

double check number 2

OpenStudy (anonymous):

and #3

OpenStudy (anonymous):

Can you explain to me how to do them?

OpenStudy (anonymous):

-(-3) = ?

OpenStudy (cwrw238):

the third question can be soved by writing the equation in the form y = mx + b where m is the slope and c - y-intercept

OpenStudy (cwrw238):

* and b = y-intercept

OpenStudy (anonymous):

Okay thank you!

OpenStudy (anonymous):

Could you help me on another?

OpenStudy (cwrw238):

2x + 3y = 18 3y = -2x + 18 divide both sides by 3: y = (-2/3) x + 6 slope = -2/3 and y-intercept = (0,6)

OpenStudy (cwrw238):

no gotta go but look again at question 2 and what completeidiot has written

OpenStudy (anonymous):

okay thanks